the sum of two numbers is -63. the first number minus the second is -41. find the numbers.

Answer by checkley77(12569)

x+y=-63
x-y=-41 add them.
--------------
2x=-104
x=-104/2
x=-52 ans for one number.
-52+y=-63
y=-63+52
y=11 ans for the other number.
Proof:
-52-11=-41
-41=-41